# LEGO: Learning and Graph-Optimized Modular Tracker for Online Multi-Object Tracking with Point Clouds

### v4

@@ III-D Feature Extraction Module
III-D1 3D Point Cloud Feature Extractor
To begin the process, the 3D point cloud is extracted from within the 3D bounding boxes using information from both the detected and predicted states. The extracted point cloud undergoes spatial adjustments using a transformation network (T-Net) before being passed through the PointNet architecture. The PointNet architecture plays a key role in generating a feature representation of the 3D point cloud data, resulting in a feature map denoted as $\bm{F}_{3d}=\left(\bm{f}_{3d_{1}},\bm{f}_{3d_{2}},\bm{f}_{3d_{3}},\dots,\bm{f}_{3d_{n}}\right)$. Each element in the feature map represents the feature vector of the object present within the 3D bounding boxes.
III-D2 Feature Embedding
Sequentially, the 3D point cloud feature map is fed into the MLP-based embedding network, to extract embedded features. This process can be described as
$\bm{F}=\left(MLP\left(\bm{F}_{3d}\right)\right),$ (4)
where $\bm{Z}^{0}$ denotes the output feature maps of the feature extraction module.
@@ III-E AMGN Score Calculation Module
Utilizing the feature map extracted from the feature extraction module, a AMGN score matrix, denoted as $\bm{B}$, is calculated through the AMGN score calculation module. Similar to [46], the purpose of employing such matrix is to extract the similarities between different detections and predicted tracks. This calculation takes into consideration the relationships between point cloud voxels of various objects. We first construct the graph. Most of the recent works which applied GNN for 3D MOT, e.g. PTP [62], Rethinking3DMOT [43] , 3Dmotformer [63], etc, use the fully connected graph with dimension of $\left(J+U\right)\times\left(J+U\right)$. However, such graph entails redundant computations due to the simultaneous inclusion of detection and prediction nodes. To this end, we follow GNN3DMOT [50] to use the prior knowledge that the matching should only happen across frames that we construct a bipartite graph. The structure of the graph is represented as a matrix of dimensions $J\times U$, where $J$ represents the number of detection nodes and $U$ denotes the number of prediction nodes. The bipartite graph mitigate this computational redundancy, effectively reducing the dimension of the association score matrix to $J\times U$. Consequently, the association between predicted and detected objects is framed as a maximum-weighted bipartite matching problem, wherein each edge in the graph corresponds to an association score between two distinct nodes. Formally, the bipartite graph is represented as $\bm{G}=\langle\bm{N},\bm{E}\rangle$, with $\bm{N}$ consisting of two node sets, $\bm{N}_{det}$ (the set of nodes representing detected objects) and $\bm{N}_{pre}$ (the set of nodes representing predicted objects). The Cardinalities of these sets are $\left|\bm{N}_{det}\right|$ = $J$ and $\left|\mathbf{N}_{pre}\right|=U$, respectively. The set of edges, $\bm{E}$, captures the associations between nodes.
@@
Figure 4: AMGN score calculation module. In the initial stage, a bipartite graph is fed into the multi-head self-attention (MHSA) block along with an LSTM block to extract features. This data, together with the initialization of the learnable adjacency matrix, are fed into the AMGN module. The AMGN module performs node message passing and updating using the self-attention coefficients derived from the learnable adjacency matrix. Subsequently, the learnable adjacency matrix is updated with the new feature representation obtained after the message passing and updating process. The new feature representation and updated learnable adjacency matrix are then fed into the next AMGN module. The final output is the AMGN score matrix $\bm{B}$, which is the latest learnable adjacency matrix from the last AMGN module.
The method in [68] directly applies polynomials to update the feature information on the node through message passing. This could result in redundant information for each node since different nodes should contribute different useful information instead of having the same contribution for message passing. To overcome the limitations in [68], GNN3DMOT [50] uses Graph Attention Network (GAT) [69] to realize the message passing and updating. However, GAT processes the message passing and updating node by node, ignoring the information of entire graph structure. This could miss important information in the constructed graph. To this end, different from the approach directly using self-attention like GAT, we propose the AMGN module, which introduces a learnable adjacency matrix into the calculation of self-attention during the message passing and updating process. The learnable adjacency matrix which contains the structure information of the graph, is updated iteratively as part of the training process, allowing the model to continuously refine its understanding of the graph structure and adapt to the evolving feature representations of the nodes through the layers and learn the strength of associations between any pair of nodes, ensuring that the attention mechanism considers both the node features and the graph’s structural information with the earnable adjacency matrix which reflect the structure of the whole graph. To use the AMGN module, we first use a multi-head self-attention (MHSA) encoder to simply extract the feature representation of each node. Considering the input is point cloud in two sequence frames, we also design an embedding network with LSTM and several fully connected layers to further extract the embedding representation of features as shown:
$\mathbf{Z}^{0}=LSTM(MHSA(\mathbf{F})).$ (5)
Then, AMGN modules are used to manage message passing and updating within the bipartite graph to get the final AMGN score matrix as shown in Fig. 4. We totally have K AMGN modules cross together and we initialize the learnable adjacency matrix of the graph as $\bm{B}^{0}$, its element $B^{0}_{j,u}$ represents the association score between nodes $j$ in the detected objects node set and nodes $u$ in the predicted objects set and the feature representation of each node is initialized by the Eq .5 and $\bm{Z^{0}}=(\bm{z}^{0}_{j},\dots,\bm{z}^{0}_{J},\bm{z}^{0}_{u},\dots,\bm{z}^{0}_{U})$. We have used total $K$ AMGN modules and suppose vector $\overrightarrow{\bm{W\odot Z}^{k-1}}$ is a vector that contains feature of all node in $(k-1)$-th AMGN module and the $u$-th node features $w_{u}z^{k-1}_{u}$ and $j$-th node feature $w_{j}z^{k-1}_{j}\in\overrightarrow{\bm{W\odot Z}^{k-1}}$. We also set $\overrightarrow{\bm{W\odot Z}^{k}}$ in $k$-th AMGN module. In our module, taking $k$-th AMGN module as an example, following GAT, the coefficient which is used for message passing and updating between node $j$ and node $u$ is calculated as:
$e^{k}_{j,u}=LeakyReLU(a^{T}[w_{u}z^{k-1}_{u}||w_{j}z^{k-1}_{j}]),$ (6)
where $a$ is the learnable weight vector, ||||| | denotes concatenation, and $\mathbf{W}$ is a learnable weight matrix. Different from GAT, which directly normalizes $e^{k}_{j,u}$ and does message passing and updating of node features, we modify the attention coefficients by combining them with the updated learnable adjacency matrix $\bm{B}^{k-1}$ from previous AMGN module as shown:
@@
where $B^{k-1}_{j,u}$ is the association score between node $j$ and node $u$. Then we normalize the modified attention coefficients with the softmax function as:
$\alpha^{k}_{j,u}=\frac{exp(e^{’k}_{j,u})}{\sum exp(e^{’k}_{j,u})},$ (8)
then we perform message passing and updating of node features as shown:
$z^{’k}_{j}=LeakyReLU(\sum a_{j,u}w_{u}{z^{k}_{u}}),$ (9)
we do the same for the $u$-th node to get the updated feature representation of node $u$ as $z^{’}_{u}$. Finally, we update value in the learnable adjacency matrix $\bm{B}^{k}$ as shown:
$B^{k}_{j,u}=LeakyReLU(\mathbf{b}^{T}[z^{’k}_{j}||z^{’k}_{u}]),$ (10)
where $\mathbf{b}^{T}$ represents the learnable weight vector at the $k$-th module. The output of the $k$-th AMGN module, which includes the learnable adjacency matrix $\bm{B}^{k}$ and the updated vector $\overrightarrow{\bm{W\odot Z}^{{}^{\prime}k}}$, serves as the input to the subsequent $(k+1)$-th AMGN module. Ultimately, the AMGN score matrix $\bm{B}$ is defined as $\bm{B}^{K}$, which is the output of the final AMGN module.
@@
$L_{a}=\sum_{j}^{J}\sum_{u}^{U}-y_{j,u}\log B_{j,u}-\left(1-y_{j,u}\right)\log\left(1-B_{j,u})\right).$ (11)
This formula represents a cross-entropy loss that measures the difference between the predicted association scores ($B_{j,u}$) and the ground truth of the matching pair ($y_{j,u}$). Beware $y_{j,u}$ is a binary indicator which is either 00 or $1$ as it represents the ground truth of the match between the $u$-th predicted object and the $j$-th detected object. When there is a clear match, $y_{j,u}$ is assigned as 1, and 0 otherwise. Due to the absence of ground-truth annotations for each pair, we define a match based on object identity continuity between time frames: a pair is considered as ”matched” if the closest ground-truth box to the tracking box in the previous time frame and the closest ground-truth box to the detection box in the current time frame have the same index and their IoU with closest ground-truth boxes are more than 0.7. The value of $B_{j,u}$ is obtained from the AMGN score matrix $\bm{B}$ and ranges between 0 and 1.
@@ III-F Association Module
$\bm{Ge}=IoU\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right)+CEN\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right).$ (13)
The IoU in Eq. (13) is the cost from 3D IoU, which can be computed as
$IoU=\frac{VI\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right)}{VOL\left(\bm{Y}_{t}^{cor}\right)+VOL\left(\bm{X}_{t|t-1}\right)-VI\left(\bm{Y}_{t}^{cor}\right)},$ (14)
where VOL is the volume calculated by $w,h,l$ in the state and VI denotes the volume intersection. The CEN in Eq. (13) is the cost from the centroid (in 3D Cartesian coordinate) [44] which can be computed as
$CEN\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right)=MSE\left(\bm{p},\bm{\hat{p}}\right),$ (15)
@@
$Mo_{u,j}=w_{ang}*\left(1-\cos\left<\tilde{v},\hat{v}\right>\right)+w_{velo}*MSE\left(\tilde{v},\hat{v}\right),$ (16)
@@
III-G State Update Module
Upon acquiring the association map from the association module, the Kalman filter is utilized to update the state of the predicted state, the same as in most of the existing works. The updated mean $\bm{X}_{t}$ and covariance $\bm{P}_{t}$ are computed through the following Kalman update equations:
$\bm{X}_{t}=\bm{X}_{t|t-1}+\bm{K}_{f}\left(\bm{Y}_{t}-\bm{H}\bm{X}_{t|t-1}\right),$ (18)
$\bm{P}_{t}=\bm{P}_{t|t-1}-\bm{K}_{f}\bm{S}\bm{K}_{f}^{T},$ (19)
$\bm{S}=\bm{H}\bm{P}_{t|t-1}\bm{H^{T}}+\bm{R},$ (20)
$\bm{K}_{f}=\bm{P}_{t|t-1}\bm{H^{T}}\left(\bm{H}\bm{P}_{t|t-1}\bm{H}^{T}+\bm{R}\right)^{-1},$ (21)
where $\mathbf{H}$ denotes the observation matrix selecting out the positional component from the object state, and $\mathbf{R}$ is the measurement noise covariance matrix.
III-H Track Management Module
In addition to the matched predicted states and the detected states, we should also deal with unmatched detected states, which may arise in two different cases: 1). Objects naturally disappear from the field of vision, and 2). Objects are misdetected, possibly due to temporary occlusion by other objects or movement beyond the optimal range of the LiDAR sensor.
In order to distinguish between these two cases, a threshold $N^{t}$ is introduced. When a predicted state fails to be associated with any detected object for more than $N^{t}$ consecutive time frames, it is presumed to have naturally disappeared. In such an event, the predicted state is eliminated and is no longer tracked. Conversely, predicted states that do not meet the above criteria are retained on the presumption that the objects may have been momentarily missed by the detector and could reappear in subsequent time.
@@ IV-A Dataset and Evaluation Metrics
The popular LiDAR-based 3D detection and tracking benchmark dataset is KITTI [71]. In this work, we use the KITTI dataset as many LiDAR-based MOT methods have been evaluated within this dataset. The efficacy of our proposed LEGO modular tracker is assessed through the Higher-Order Tracking Accuracy (HOTA), defined as
[…]
$HOTA_{\alpha}=\sqrt{\frac{\sum_{c}A(c)}{FN+FP+TP}},$ (23)
@@
$AMOTA=\frac{1}{L}\sum_{r}\left(1-\frac{\sum_{t}FN_{t}^{r}+FP_{t}^{r}+IDS_{t}^{r}}{\sum_{t}GT_{t}}\right),$ (26)
where $FN_{t}^{r}$, $FP_{t}^{r}$, $IDS_{t}^{r}$ represent the number of false negatives, false positives, ID switch at a specific recall value $r$ at time $t$, $L$ is the number of recall values respectively and AMOTP is defined as
$AMOTP=\frac{1}{L}\sum_{r}\left(1-MOTP_{r}\right),$ (27)
where $MOTP_{r}$ is the value of MOTP at a specific recall ${r}$.
(a) Tracking ground truth for scene 0006
(b) LEGO tracking results for scene 0006
(c) Tracking ground truth for scene 0008
(d) LEGO tracking results for scene 0008
Figure 5: This represents a demonstration of 3D MOT results utilizing the KITTI benchmark. The first and second rows display the tracking ground truth values and our proprietary LEGO model results, specifically within scene 0006. Within this context, the detected state (after the offset correction module) in the current frame is signified by the presence of a bounding box and an associated tracking identification number. The same objects have bounding boxes in the same colour with the same identification number. Similarly, the third and fourth rows offer an exhibition of the ground truth and employ our LEGO model results, albeit within a different context, scene 0008. These rows continue to utilize the detected state (after the offset correction module), marked by a bounding box and the tracking identification, to illuminate tracking methods within this unique environment.

### v6

@@ III-D Feature Extraction Module
Specifically, the 3D point clouds within the detected and predicted 3D bounding boxes are cropped, feed into a PointNet encoder [73] to generate corresponding feature maps, denoted as $\bm{F}^{det}_{3d}=\left(\bm{f}^{det}_{3d_{1}},\bm{f}^{det}_{3d_{2}},\dots,\bm{f}^{det}_{3d_{J}}\right)$ and $\bm{F}^{pred}_{3d}=\left(\bm{f}^{pred}_{3d_{1}},\bm{f}^{pred}_{3d_{2}},\dots,\bm{f}^{pred}_{3d_{U}}\right)$. Sequentially, the 3D point cloud feature maps are fed into the MLP-based embedding network, to extract embedded features. This process can be described as
$\mathbf{F}=\mathtt{MLP}\left(\bm{F}^{det}_{3d}\right),$ (5)
$\mathbf{F^{*}}=\mathtt{MLP}\left(\bm{F}^{pred}_{3d}\right),$ (6)
where $\mathbf{F}$ and $\mathbf{F^{*}}$ denote the output feature maps of the feature extraction module.
@@ III-E AMGN Score Calculation Module
Utilizing the feature map extracted from the feature extraction module, an AMGN score matrix, denoted as $\bm{B}$, is calculated through the AMGN score calculation module. Like [47], the purpose of employing such a matrix is to measure the similarity between different detections and predictions. This section details the AMGN Score Calculation Module in three parts: graph construction, AMGN blocks, and the loss function.
III-E1 Graph Construction
Most recent works, which applied GNN for 3D MOT, e.g., PTP [67], Rethinking3DMOT [44], 3Dmotformer [68], use a fully connected graph with the dimension of $\left(J+U\right)\times\left(J+U\right)$ where $J$ and $U$ denote the number of detected nodes and the number of predicted nodes in the graph, respectively. However, such a graph entails redundant computations due to the simultaneous inclusion of detection and prediction nodes.
To this end, following the approach of GNN3DMOT [52], the prior knowledge that matching should only occur across frames is utilized, and a bipartite graph is constructed. The structure of the graph is represented as a matrix of dimensions $J\times U$. The bipartite graph mitigates this computational redundancy, effectively reducing the dimension of the association score matrix. Formally, the bipartite graph is represented as $\bm{G}=\langle\bm{N},\bm{E}\rangle$, with the nodes $\bm{N}$ are split into two distinct sets: one for detected objects $\bm{N}_{det}$ and one for predicted objects $\bm{N}_{pre}$. The set $\bm{E}$ of edges captures the associations between the nodes.
@@ III-E1 Graph Construction
Figure 3: AMGN score calculation module. In the initial stage, a bipartite graph is fed into the multi-head self-attention (MHSA) block along with an LSTM block to extract features. This data, together with the initialization of the learnable adjacency matrix, are fed into the AMGN block. The AMGN block performs node message passing and updating using the self-attention coefficients derived from the learnable adjacency matrix. Subsequently, the learnable adjacency matrix is updated with the new feature representation obtained after the message passing and updating process. The new feature representation and updated learnable adjacency matrix are then fed into the next AMGN block. The final output is the AMGN score matrix $\bm{B}$, which is the latest learnable adjacency matrix from the last AMGN block.
III-E2 AMGN Blocks
GNN3DMOT [52] leverages the Graph Attention Network (GAT) [75] to enable different contributions from individual nodes during message passing. However, GAT processes nodes independently, and thus it may fail to capture important contextual information embedded in the global graph structure.
To overcome this, the AMGN block is proposed, which introduces a learnable adjacency matrix into the self-attention calculation. This matrix contains the structural information of the graph, allowing the model to continuously refine its understanding of the graph structure, adapt to the evolving feature representations of the nodes, and learn the strength of associations between any pair of nodes. This approach ensures that the attention mechanism incorporates both the node features and the global structural information of the graph, as captured by the learnable adjacency matrix.
The module’s process begins by generating robust initial embeddings, which are then iteratively refined by K AMGN blocks. First, a Multi-Head Self-Attention (MHSA) encoder extracts preliminary features from the concatenated detected and predicted feature maps. An LSTM network then processes these features to enhance temporal coherence, generating refined initial embeddings:
$\mathbf{Z}^{0}=\mathtt{LSTM}(\mathtt{MHSA}([\mathbf{F},\mathbf{F}^{*}])),$ (7)
where $\bm{Z^{0}}$ is the bipartite graph with the embedding representation sets of detected nodes and predicted nodes that contains $\bm{z}^{0}_{0},\dots,\bm{z}^{0}_{j},\dots,\bm{z}^{0}_{J}]$ and $\bm{z}^{*0}_{0},\dots,\bm{z}^{*0}_{u},\dots,\bm{z}^{*0}_{U}]$.
After obtaining the initial embeddings, the K AMGN blocks iteratively refine these embeddings and update the learnable adjacency matrix as shown in Fig. 3. The learnable adjacency matrix is initialized as $\bm{B}^{0}$, where its element $B^{0}_{j,u}$ between the $j$-th node in the set of nodes for detected objects and the $u$-th node in the set of nodes for predicted objects.
For the k-th AMGN block, a raw attention coefficient $e^{k}_{j,u}$ between $j$-th detected node and $u$-th predicted node is first calculated as:
$e^{k}_{j,u}=\mathtt{LeakyReLU}(a^{T}[w^{*}_{u}z^{*k-1}_{u}||w_{j}z^{k-1}_{j}]),$ (8)
where $a$ is the learnable weight vector, $||$ denotes concatenation, and $w^{*}_{u},w_{j}$ are elements in a learnable weight matrix.
Different from GAT, these attention coefficients are modified by combining them with the learnable adjacency matrix $\bm{B}^{k-1}$ from the previous block as shown:
@@ III-E2 AMGN Blocks
where $B^{k-1}_{j,u}$ is the association score between $j$-th detected node and $u$-th predicted node from the prior step.
The modified coefficients are then normalized using a softmax function as:
$\alpha^{k}_{j,u}=\frac{\mathrm{exp}(e^{’k}_{j,u})}{\sum\mathrm{exp}(e^{’k}_{j,u})}.$ (10)
With these normalized attention coefficients, the node embedding is updated by aggregating neighbour information. The updated embedding for $j$-th detected node is calculated as:
$z^{’k}_{j}=\mathtt{LeakyReLU}(\sum a_{j,u}w_{u}{z^{*k}_{u}}).$ (11)
We do the same for the $u$-th predicted node to get the updated feature representation as $z^{’*k}_{u}$.
Finally, the value in the learnable adjacency matrix $\bm{B}^{k}$ is updated as shown:
$B^{k}_{j,u}=\mathtt{LeakyReLU}(\mathbf{b}^{T}[z^{’k}_{j}||z^{’*k}_{u}]),$ (12)
where $\mathbf{b}^{T}$ represents the learnable weight vector at the $k$-th module. The output of the $k$-th AMGN block serves as the input to the subsequent $(k+1)$-th AMGN block. Ultimately, the AMGN score matrix $\bm{B}$ is defined as $\bm{B}^{K}$, which is the output of the final AMGN block.
III-E3 Loss Function
@@
$L_{a}=\sum_{j}^{J}\sum_{u}^{U}\left[-y_{j,u}\log B_{j,u}-\left(1-y_{j,u}\right)\log\left(1-B_{j,u}\right)\right].$ (13)
This formula represents a cross-entropy loss that measures the difference between the predicted association scores ($B_{j,u}$) and the ground truth of the matching pair ($y_{j,u}$). Beware that $y_{j,u}$ is a binary indicator, which is either 0 or $1$, as it represents the ground truth of the match between the $u$-th predicted node and the $j$-th detected node. When there is a clear match, $y_{j,u}$ is assigned as 1, and 0 otherwise. Due to the absence of ground-truth annotations for each pair, a match is determined based on the continuity of object identity across time frames. In the match, a pair is considered as matched if the closest ground-truth box to the tracking box in the previous time frame and the closest ground-truth box to the detection box in the current time frame have the same index and their IoU with the closest ground-truth box is more than 0.7. The value of $B_{j,u}$ is obtained from the AMGN score matrix $\bm{B}$ and ranges between 0 and 1.
@@ III-F Association Module
$\bm{Ge}=\mathtt{IoU}\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right)+\mathtt{CEN}\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right).$ (15)
The $\mathtt{IoU}$ in Eq. (15) is the cost from 3D IoU, which can be computed as
$\mathtt{IoU}=\frac{\mathtt{VI}\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right)}{\mathtt{VOL}\left(\bm{Y}_{t}^{cor}\right)+\mathtt{VOL}\left(\bm{X}_{t|t-1}\right)-\mathtt{VI}\left(\bm{Y}_{t}^{cor},\bm{X}_{t|t-1}\right)},$ (16)
where $\mathtt{VOL}$ is the volume calculated by $w,h,l$ in the state and $\mathtt{VI}$ denotes the volume intersection. The $\mathtt{CEN}$ in Eq. (15) is the cost from the centroid (in 3D Cartesian coordinate) [45] which can be computed as
[…]
@@
[…]
@@ IV Experiments and Analysis
    The metrics are reported by using CasA [16] as 3D object detector.
@@
TABLE IV: 3D MOT tracking results of the proposed method and other state-of-the-art trackers on Waymo [80] validation dataset.
[…]
SimTrack (ICCV 2021)+ [25] L 53.1 49.6 17.4 17.4
3DMODT (IEEE ICRA 2023)+ [81] L 55.9 51.2 18.9 18.9
SimpleTrack (ECCV Workshop 2022)* [28] L $N/A$ 56.9 $N/A$ $N/A$
Waymo LEGO (Ours)* L 58.30 58.30 19.44 19.44
PC3T (IEEE T-ITS 2022)** [45] L 61.3 61.3 $N/A$ $N/A$
Fast-Poly (IEEE RA-L 2024)** [82] L 62.3 62.3 $N/A$ $N/A$
LEGO (Ours)** L 63.1 63.1 23.6 23.6
    The metrics are reported without using pre-trained 3D object detector, but a joint detection and tracking network delicately for this task.
    The metrics are reported by using CenterPoint [15] as 3D object detector.
    The metrics are reported by using CasA [16] as 3D object detector.
@@ IV-A Dataset and Evaluation Metrics
In this work, we use the KITTI dataset [78] and Waymo dataset [79] as many LiDAR-based MOT methods have been evaluated within these datasets. The efficacy of our proposed LEGO modular tracker is assessed through the Higher-Order Tracking Accuracy (HOTA) [77], defined as
$HOTA=\int_{0}^{1}HOTA_{\alpha}d\alpha\approx\frac{1}{19}\sum_{\alpha}HOTA_{\alpha},$ (20)
$HOTA_{\alpha}=\sqrt{\frac{\sum_{c}\mathcal{A}(c)}{FN+FP+TP}},$ (21)
@@
$AMOTA=\frac{1}{M}\sum_{r}\left(1-\frac{\sum_{t}FN_{t}^{r}+FP_{t}^{r}+IDS_{t}^{r}}{\sum_{t}GT_{t}}\right),$ (24)
where $FN_{t}^{r}$, $FP_{t}^{r}$, $IDS_{t}^{r}$ represent the number of false negatives, false positives, ID switch at a specific recall value $r$ at time $t$, $M$ is the number of recall values respectively and AMOTP is defined as
$AMOTP=\frac{1}{M}\sum_{r}\left(1-MOTP_{r}\right),$ (25)
where $MOTP_{r}$ is the value of MOTP at a specific recall r{r}.
(a) Tracking ground truth for scene 0008
(b) PC3T tracking results for scene 0008
(c) LEGO tracking results for scene 0008
Figure 4: Visualization of 3D MOT results on the KITTI benchmark (scene 0008). The first row shows the ground truth tracking annotations across three consecutive frames. Each detected object in the current frame is represented by a bounding box along with its unique tracking ID. Objects are consistently colored according to their tracking IDS across frames. The second and third rows illustrate the tracking results obtained by PC3T and our proposed LEGO model, respectively, on the same frames. Please zoom in to see the details.